OpenStudy (anonymous):

Evaulate the limits lim ((1/s)-(1/6))/(s-6) s->6

OpenStudy (anonymous):

Typing it in latex makes it a bit unclear, so I'll do my best typing it normally: lim s→6 [1/s - 1/6] / (s - 6) The numerator's fractions can be combined to get (6 - s)/(6s). And this over (s - 6), so you have lim s→6 [(6 - s)/(6s)] / (s - 6) Note that s - 6 = -(6 - s), so you have (-1) lim s→6 [(6 - s)/(6s)] / (6 - s) You can probably see where this is going.

OpenStudy (anonymous):

mmmmm

OpenStudy (whpalmer4):

I'm not sure why it should be any less clear when formatted in LaTeX. It is a pain to type, though! \[\lim_{s\rightarrow 6 } \frac{(\dfrac{1}{s}-\dfrac{1}{6})}{s-6}\]Establish a common denominator for the numerator fractions \[\lim_{s\rightarrow 6 } \frac{(\dfrac{1*6}{s*6}-\dfrac{1*s}{6*s})}{s-6} = \lim_{s \rightarrow 6} [\frac{6-s}{6s}*\frac{1}{s-6}] = \lim_{s \rightarrow 6} [\frac{-1(s-6)}{6s}*\frac{1}{s-6} ] =\]and hopefully you can do the rest!

OpenStudy (anonymous):

Yeah, I didn't like how small the s turns out under the limit.